# ShiftAddViT: Mixture of Multiplication Primitives Towards Efficient Vision Transformers

### v4

@@ 4.2 ShiftAddViT: Mixture of Experts Framework
Hypothesis. We hypothesize there are important yet sensitive input tokens that necessitate the more powerful networks otherwise suffer from accuracy drops. Those tokens are likely to contain object information that directly correlates with our task objectives as validated by bolya2022token ; yu2022mia ; rao2021dynamicvit ; yin2022avit . While the left tokens are less sensitive and can be fully represented even using cheaper Shift layers. Such an input-adaptive nature of ViTs calls for a framework featuring a mixture of multiplication primitives.
@@
  • Dynamic Input Allocation. The input allocation is determined dynamically during runtime, and the trainable router within the MoE layers automatically learns this dispatch assignment as the gradients decay. We know the allocation only when the model is executed and we receive the router outputs. Therefore, the shape of expert input and corresponding indexes are dynamically changed. This can be handled by PyTorch with dynamic graphs while TVM expects static input shape. To this end, we follow and leverage the compiler support for dynamism as proposed in Nimble shen2021nimble on top of the Apache TVM to handle the dynamic input allocation.
  • Parallel Computing. Different experts are expected to run in parallel, this can be supported by several customized distributed training frameworks integrated with PyTorch, e.g., FasterMoE he2022fastermoe , and DeepSpeed rajbhandari2022deepspeed . However, it remains nontrivial to support this using TVM. One option to simulate is to perform modularized optimization to mimic parallel computing, in particular, we optimize each expert separately and measure their latency, the maximum latency among all experts will be recorded and regarded as the latency of this MoE layer, and the aggregated total of the time spent for each layer is the final reported modularized model latency. To avoid any potential confusion between real-measured wall-clock time, i.e., no parallelism assumed, and simulated modularized latency, i.e., ideal parallelism assumed, we reported both for models containing MoE layers in Sec. 5 to offer more insights into the algorithm’s feasibility and cost implications.
Latency-aware Load-balancing Loss. The key to our MoE framework is the design of a routing function to balance all experts towards higher accuracy and lower latency. Previous solutions fedus2022switch ; he2021fastmoe use homogeneous experts and treat them equally. While in our MoE framework, the divergence and heterogeneity between powerful yet slow Mult. and fast yet less powerful Shift experts incur one unique question: How to orchestrate the workload of each expert to reduce the synchronization time? Our answer is a latency-aware load-balancing loss, which ensures (1) all experts receive the expected weighted sums of gate values; and (2) all experts are assigned the expected number of input tokens. Those two conditions are enforced by adopting importance and load balancing loss as defined below:
@@
where SCV denotes the squared coefficient of variation of given distributions over experts (also used in riquelme2021scaling ; shazeer2017outrageously ); $\alpha_{i}$ refers to the latency-aware coefficient of $i$-th expert and is defined by $\nicefrac{{(\texttt{Lat}_{i})}}{{(\sum_{j}\texttt{Lat}_{j})}}$ because that the expected assignments are inversely proportional to runtime latency of $i$-th expert, $\texttt{Lat}_{i}$; $q_{i}(\mathbf{x})$ is the probability that the gate value of $i$-th expert outweighs the others, i.e., top1 expert, and is given by $\mathbf{P}\left(p_{i}(\mathbf{x})+\epsilon\geq p_{j}(\mathbf{x}),\forall j\neq i\right)$. Note that this probability is discrete, we use a noise proxy $\epsilon$ following shazeer2017outrageously ; he2021fastmoe to make it differentiable. The above latency-aware importance loss and load-balanced loss help to balance gate values and workload assignments, respectively, and can be integrated with classification loss $\mathcal{L}_{\texttt{CLS}}$ as the total loss function $\mathcal{L}(\mathbf{X})=\mathcal{L}_{\texttt{CLS}}(\mathbf{X})+\lambda\cdot(\mathcal{L}_{\texttt{IMP}}(\mathbf{X})+\mathcal{L}_{\texttt{LOAD}}(\mathbf{X}))$ for training ViTs and gates simultaneously. $\lambda$ is set as 0.01 for all experiments.
@@ 5.1 Experiment Settings
Models and Datasets. Tasks and Datasets. We consider two representative 2D and 3D Transformer-based vision tasks to demonstrate the superiority of the proposed ShiftAddViT, including 2D image classification on ImageNet dataset deng2009imagenet with 1.2 million training and 50K validation images and 3D novel view synthesis (NVS) task on Local Light Field Fusion (LLFF) dataset mildenhall2019local with eight scenes. Models. For the classification task, we consider PVTv1 PVT_ICCV , PVTv2 wang2022pvt , and DeiT DeiT . For the NVS task, we consider Transformer-based GNT GNT with View- and Ray-Transformer models.
Training and Inference Details. For the classification task, we follow Ecoformer liu2022ecoformer to initialize the pre-trained ViTs with Multihead Self-Attention (MSA) weights, based on which we apply our reparameterization a two-stage finetuning: (1) convert MSA to linear attention you2023castling and reparameterize all MatMuls with add layers with 100 epoch finetuning, and (2) reparameterize MLPs or linear layers with shift or MoE layers after finetuning another 100 epoch. Note that we follow PVTv2 wang2022pvt and Ecoformer to keep the last stage as MSA for fair comparisons. For the NVS task, we still follow the two-stage finetuning but do not convert MSA weights to linear attention to maintain the accuracy. All experiments are run on a server with eight RTX A5000 GPUs with each having 24GB GPU memory.
Baselines and Evaluation Metrics. Baselines. For the classification task, we reparameterize and compare our ShiftAddViT with PVTv1 PVT_ICCV , PVTv2 wang2022pvt , Ecoformer liu2022ecoformer , and their MSA variants. For the NVS task, we reparameterize on top of GNT GNT and compare with both vanilla NeRF mildenhall2021nerf and GNT GNT . Evaluation Metrics. For the classification task, we evaluate the ShiftAddViT and baselines in terms of accuracy, GPU latency and throughput measured on an RTX 3090 GPU. For the NVS task, we evaluate all models in terms of PSNR, SSIM wang2004image , and LPIPS zhang2018unreasonable . We also measure and report the energy consumption of all the above models based on an Eyeriss-like hardware accelerator chen2016eyeriss ; zhao2020dnn , which calculates not only computational but also data movement energy.
@@ 5.2 ShiftAddViT over SOTA Baselines on 2D Tasks
Table 4: Comparisons between ShiftAddViT and baselines. We show the breakdown analysis of ShiftAddViT with two kinds of $\mathbf{Q/K}$ quantization. The throughput or latency is measured with a batch size of 32 or 1, where ${}^{\dagger}$ denotes that the numbers are measured and reported after optimizing ShiftAddViT using TVM. We report both real-device and modularized latency for models with MoE.
  Methods Linear Attn Add Shift MoE PVTv2-B0 wang2022pvt PVTv1-T PVT_ICCV
@@
Ecoformer liu2022ecoformer 70.44 7.82 1348 NaN 7.43 1021
ShiftAddViT (with KSH liu2022ecoformer or Quant. hubara2017quantized to binarize Q/K) 71.19 6.13 2066 75.50 5.78 1640
70.95 ${}^{\dagger}$ ${}^{\dagger}$ 75.20 ${}^{\dagger}$ ${}^{\dagger}$
✓ (Attn) 70.53 ${}^{\dagger}$ ${}^{\dagger}$ 74.77 ${}^{\dagger}$ ${}^{\dagger}$
✓ (Attn) ✓ (MLPs) 70.16 ${}^{\dagger}$$*{}^{*}$ N/A 74.44 ${}^{\dagger}$$*{}^{*}$ N/A
✓ (Both) 70.38 ${}^{\dagger}$$*{}^{*}$ N/A 74.73 ${}^{\dagger}$$*{}^{*}$ N/A
@@
71.04 ${}^{\dagger}$ ${}^{\dagger}$ 75.18 ${}^{\dagger}$ ${}^{\dagger}$
✓ (Both) 68.57 ${}^{\dagger}$ ${}^{\dagger}$ 73.47 ${}^{\dagger}$ ${}^{\dagger}$
✓ (Both) 70.59 ${}^{\dagger}$$*{}^{*}$ N/A 74.93 ${}^{\dagger}$$*{}^{*}$ N/A
@@
Table 5: Comparisons between ShiftAddViT and baselines. We apply our Shift&Add techniques on top of SOTA Transformer-based NeRF model, GNT GNT , and report averaged PSNR ($\uparrow$), SSIM ($\uparrow$) and LPIPS ($\downarrow$) on the LLFF dataset mildenhall2019local across eight scenes. In addition, we also show the results on two representative scenes, Orchids, and Flower, and the rendering latency and energy measured on an Eyeriss-like accelerator. More results for other scenes are available in the Appendix C.
@@
Training Wall-clock Time. To offer more insights into the training cost, we collect the training wall-clock time, which is 21% $\sim$ 25% less than training from scratch as the original Transformer model and more than 50$\times$ less than the previous ShiftAddNet you2020shiftaddnet did on GPUs. For example, training the PVTv1-Tiny model from scratch takes 62 hours while our finetuning only needs 46 hours.
@@ 5.3 ShiftAddViT over SOTA Baselines on 3D Tasks
We also extend our methods to 3D NVS tasks and built ShiftAddViT on top of Transformer-based GNT models GNT . We test and compare the PSNR, SSIM, LPIPS, latency, and energy of both ShiftAddViT and baselines, i.e., vanilla NeRF mildenhall2021nerf and GNT GNT , on the LLFF dataset across eight scenes. As shown in Tab. 5, our ShiftAddViT consistently leads to better accuracy-efficiency tradeoffs, achieving 22.3%/50.4% latency reductions and 20.8%/54.3% energy savings under comparable or even better generation quality ($\uparrow$0.55$\downarrow$0.19 averaged PSNR across eight scenes), as compared to NeRF and GNT baselines, respectively. Note that GNT costs more than NeRF because of an increased number of layers. In particular, ShiftAddViT even achieves better generation quality (up to $\uparrow$0.88 PSNR) than GNT on some representation scenes, e.g., Orchid and Flower, as highlighted in Tab. 5, where the first, second, and third ranking performance are noted with corresponding colors. The full results of eight scenes are supplied in the Appendix C. Note that we only finetune ShiftAddViT for 140K steps on top of pre-trained GNT models instead of training up to 840K steps from scratch.
@@ 5.4 Ablation Studies of ShiftAddViT
  Methods Linear Attn Add Shift MoE PVTv2-B1 wang2022pvt PVTv2-B2 wang2022pvt
@@
PVT wang2022pvt 78.70 6.75 1512 82.00 12.20 921
Ecoformer liu2022ecoformer 78.38 8.02 874 81.28 15.43 483
ShiftAddViT (with KSH liu2022ecoformer or Quant. hubara2017quantized to binarize Q/K) 78.80 6.36 1373 81.40 9.21 796
78.70 ${}^{\dagger}$ ${}^{\dagger}$ 81.31 ${}^{\dagger}$ ${}^{\dagger}$
✓ (Attn) ✓ (MLPs) 78.20 ${}^{\dagger}$$*{}^{*}$ N/A 81.06 ${}^{\dagger}$$*{}^{*}$ N/A
✓ (Both) 78.38 ${}^{\dagger}$$*{}^{*}$ N/A 81.32 ${}^{\dagger}$$*{}^{*}$ N/A
@@
78.70 ${}^{\dagger}$ ${}^{\dagger}$ 81.33 ${}^{\dagger}$ ${}^{\dagger}$
✓ (Both) 77.55 ${}^{\dagger}$ ${}^{\dagger}$ 79.97 ${}^{\dagger}$ ${}^{\dagger}$
✓ (Both) 78.49 ${}^{\dagger}$$*{}^{*}$ N/A 81.01 ${}^{\dagger}$$*{}^{*}$ N/A
Performance Breakdown Analysis. To investigate how each of the proposed techniques contributes to the final performance, we conduct ablation studies among various kinds of ShiftAddViT variants on PVTv1-T, PVTv2-B0/B1/B2 models to gradually break down and show the advantage of each component. As shown in Tab. 4 and 6, we have three general observations: (1) ShiftAddViT is robust to the binarization method of $\mathbf{Q/K}$, e.g., it is compatible with either KSH liu2022ecoformer or vanilla binarization hubara2017quantized , achieving on average 3.03$\times$ latency reductions than original PVT under comparable accuracies (±$\pm 0.5\%$), and both 3.29$\times$/1.32$\times$ latency/energy reductions and 0.04% $\sim$ 0.20% accuracy improvements over Ecoformer liu2022ecoformer ; (2) vanilla binarization methods work better than KSH in our ShiftAddViT framework in terms of both accuracy ($\downarrow$0.05$\sim$ $\uparrow$0.21%) and efficiency (on average $\downarrow$5.9latency reductions and $\uparrow$5.8throughput boosts). Moreover, KSH requires that $\mathbf{Q}$ and $\mathbf{K}$ are identical while vanilla binarization hubara2017quantized does not have such a limitation. That explains why applying vanilla quantization to the linear attention layers results in an accuracy improvement of on average 0.15% as compared to using KSH. Also note that our adopting linear attention following you2023castling works better ($\uparrow$0.29%) than PVT PVT_ICCV ; wang2022pvt ; (3) Replacing linear layers in attention modules or MLPs with Shift layers lead to an accuracy drop while our proposed MoE can help to compensate for it. Specifically, adopting Shift layers leads to on average 1.67% accuracy drop while MoE instead improves on average 1.37% accuracy. However, MoE hurts the efficiency of both the PVT baseline and our ShiftAddViT without customized system acceleration, e.g., PVT+MoE increases 94% latency and reduces 51% throughput than PVT on GPUs, due to limited parallelism supports, especially for TVM. We report the modularized latency by separately optimizing each expert to demonstrate the potential of MoE’s double-winning accuracy ($\uparrow$0.94$\sim$ $\uparrow$2.02%) and efficiency ($\downarrow$15.4$\sim$ $\uparrow$42.7%).
@@
Figure 4: MLP/Linear latency speedups using shifts, where inputs are of shape $B\times K\times M$, weights are of shape $K\times N$. All dimensions are set w.r.t. the real dimensions in PVTs PVT_ICCV .
Figure 4: MLP/Linear latency speedups using shifts, where inputs are of shape $B\times K\times M$, weights are of shape $K\times N$. All dimensions are set w.r.t. the real dimensions in PVTs PVT_ICCV .
@@
Speedups of Shifts and Adds. Apart from the overall comparison and energy reduction profiling, we also test the GPU speedups of our customized Shift and Add kernels, as shown in Fig. 5 and 5, respectively. We can see that both customized kernels achieve faster speeds than PyTorch and TVM baselines. For example, our MatAdds achieve on average 7.54$\times$/1.51$\times$ speedups than PyTorch and TVM MatMuls, respectively. Our MatShifts achieve on average 2.35$\times$/3.07$\times$/1.16$\times$ speedups than PyTorch FakeShifts elhoushi2021deepshift , TVM FakeShifts, and TVM MatMuls, respectively. More comparisons and speedup analysis under larger batch sizes are supplied in the Appendix A. Note that in those comparisons, we use the default einsum operator for the PyTorch MatMuls, floating-point multiplication with power-of-twos for FakeShift, and implement the MatAdds and MatShifts using TVM chen2018tvm ourselves. We compare those customized operators with their multiplication counterparts in the same setting for fair comparisons. The speedup of MatAdds is because we replace multiplication-and-accumulation with accumulation only. The speedup of MatShifts is mainly attributed to the bit reductions (INT32 and INT8 for inputs and shift signs/weights) and thus data movement reductions instead of computations which are almost fully hidden behind data movements.

### v6

@@ 4.2 ShiftAddViT: Mixture of Experts Framework
Hypothesis. We hypothesize there are important yet sensitive input tokens that necessitate the more powerful networks otherwise suffer from accuracy drops. Those tokens are likely to contain object information that directly correlates with our task objectives as validated by [3, 74, 45, 68]. While the left tokens are less sensitive and can be fully represented even using cheaper Shift layers. Such an input-adaptive nature of ViTs calls for a framework featuring a mixture of multiplication primitives.
@@
  • Dynamic Input Allocation. The input allocation is determined dynamically during runtime, and the trainable router within the MoE layers automatically learns this dispatch assignment as the gradients decay. We know the allocation only when the model is executed and we receive the router outputs. Therefore, the shape of expert input and corresponding indexes are dynamically changed. This can be handled by PyTorch with dynamic graphs while TVM expects static input shape. To this end, we follow and leverage the compiler support for dynamism as proposed in Nimble [49] on top of the Apache TVM to handle the dynamic input allocation.
  • Parallel Computing. Different experts are expected to run in parallel, this can be supported by several customized distributed training frameworks integrated with PyTorch, e.g., FasterMoE [24], and DeepSpeed [44]. However, it remains nontrivial to support this using TVM. One option to simulate is to perform modularized optimization to mimic parallel computing, in particular, we optimize each expert separately and measure their latency, the maximum latency among all experts will be recorded and regarded as the latency of this MoE layer, and the aggregated total of the time spent for each layer is the final reported modularized model latency. To avoid any potential confusion between real-measured wall-clock time, i.e., no parallelism assumed, and simulated modularized latency, i.e., ideal parallelism assumed, we reported both for models containing MoE layers in Sec. 5 to offer more insights into the algorithm’s feasibility and cost implications.
Latency-aware Load-balancing Loss. The key to our MoE framework is the design of a routing function to balance all experts towards higher accuracy and lower latency. Previous solutions [20, 23] use homogeneous experts and treat them equally. While in our MoE framework, the divergence and heterogeneity between powerful yet slow Mult. and fast yet less powerful Shift experts incur one unique question: How to orchestrate the workload of each expert to reduce the synchronization time? Our answer is a latency-aware load-balancing loss, which ensures (1) all experts receive the expected weighted sums of gate values; and (2) all experts are assigned the expected number of input tokens. Those two conditions are enforced by adopting importance and load balancing loss as defined below:
@@
where SCV denotes the squared coefficient of variation of given distributions over experts (also used in [46, 48]); $\alpha_{i}$ refers to the latency-aware coefficient of $i$-th expert and is defined by $\nicefrac{{(\texttt{Lat}_{i})}}{{(\sum_{j}\texttt{Lat}_{j})}}$ because that the expected assignments are inversely proportional to runtime latency of $i$-th expert, $\texttt{Lat}_{i}$; $q_{i}(\mathbf{x})$ is the probability that the gate value of $i$-th expert outweighs the others, i.e., top1 expert, and is given by $\mathbf{P}\left(p_{i}(\mathbf{x})+\epsilon\geq p_{j}(\mathbf{x}),\forall j\neq i\right)$. Note that this probability is discrete, we use a noise proxy $\epsilon$ following [48, 23] to make it differentiable. The above latency-aware importance loss and load-balanced loss help to balance gate values and workload assignments, respectively, and can be integrated with classification loss $\mathcal{L}_{\texttt{CLS}}$ as the total loss function $\mathcal{L}(\mathbf{X})=\mathcal{L}_{\texttt{CLS}}(\mathbf{X})+\lambda\cdot(\mathcal{L}_{\texttt{IMP}}(\mathbf{X})+\mathcal{L}_{\texttt{LOAD}}(\mathbf{X}))$ for training ViTs and gates simultaneously. $\lambda$ is set as 0.01 for all experiments.
@@ 5.1 Experiment Settings
Models and Datasets. Tasks and Datasets. We consider two representative 2D and 3D Transformer-based vision tasks to demonstrate the superiority of the proposed ShiftAddViT, including 2D image classification on ImageNet dataset [15] with 1.2 million training and 50K validation images and 3D novel view synthesis (NVS) task on Local Light Field Fusion (LLFF) dataset [40] with eight scenes. Models. For the classification task, we consider PVTv1 [60], PVTv2 [61], and DeiT [55]. For the NVS task, we consider Transformer-based GNT [53] with View- and Ray-Transformer models.
Training and Inference Details. For the classification task, we follow Ecoformer [34] to initialize the pre-trained ViTs with Multihead Self-Attention (MSA) weights, based on which we apply our reparameterization a two-stage finetuning: (1) convert MSA to linear attention [73] and reparameterize all MatMuls with add layers with 100 epoch finetuning, and (2) reparameterize MLPs or linear layers with shift or MoE layers after finetuning another 100 epoch. Note that we follow PVTv2 [61] and Ecoformer to keep the last stage as MSA for fair comparisons. For the NVS task, we still follow the two-stage finetuning but do not convert MSA weights to linear attention to maintain the accuracy. All experiments are run on a server with eight RTX A5000 GPUs with each having 24GB GPU memory.
Baselines and Evaluation Metrics. Baselines. For the classification task, we reparameterize and compare our ShiftAddViT with PVTv1 [60], PVTv2 [61], Ecoformer [34], and their MSA variants. For the NVS task, we reparameterize on top of GNT [53] and compare with both vanilla NeRF [41] and GNT [53]. Evaluation Metrics. For the classification task, we evaluate the ShiftAddViT and baselines in terms of accuracy, GPU latency and throughput measured on an RTX 3090 GPU. For the NVS task, we evaluate all models in terms of PSNR, SSIM [63], and LPIPS [76]. We also measure and report the energy consumption of all the above models based on an Eyeriss-like hardware accelerator [12, 77], which calculates not only computational but also data movement energy.
@@ 5.2 ShiftAddViT over SOTA Baselines on 2D Tasks
Table 4: Comparisons between ShiftAddViT and baselines. We show the breakdown analysis of ShiftAddViT with two kinds of $\mathbf{Q/K}$ quantization. The throughput or latency is measured with a batch size of 32 or 1, where denotes that the numbers are measured and reported after optimizing ShiftAddViT using TVM. We report both real-device and modularized latency for models with MoE.
     Methods Linear Attn Add Shift MoE         PVTv2-B0 [61]         PVTv1-T [60]
@@
Ecoformer [34] 70.44 7.82 1348 NaN 7.43 1021
ShiftAddViT (with KSH [34] or Quant. [27] to binarize Q/K) 71.19 6.13 2066 75.50 5.78 1640
70.95 1.07 2530 75.20 1.42 1683
✓ (Attn) 70.53 1.04 2447 74.77 1.39 1647
✓ (Attn) ✓ (MLPs) 70.16 1.39/1.11 N/A 74.44 1.91/1.21 N/A
✓ (Both) 70.38 1.59/1.20 N/A 74.73 2.12/1.21 N/A
@@
71.04 1.00 2613 75.18 1.20 1907
✓ (Both) 68.57 0.97 2736 73.47 1.18 1820
✓ (Both) 70.59 1.51/1.12 N/A 74.93 1.97/1.02 N/A
@@
Table 5: Comparisons between ShiftAddViT and baselines. We apply our Shift&Add techniques on top of SOTA Transformer-based NeRF model, GNT [53], and report averaged PSNR ($\uparrow$), SSIM ($\uparrow$) and LPIPS ($\downarrow$) on the LLFF dataset [40] across eight scenes. In addition, we also show the results on two representative scenes, Orchids, and Flower, and the rendering latency and energy measured on an Eyeriss-like accelerator. More results for other scenes are available in the Appendix C.
@@
Training Wall-clock Time. To offer more insights into the training cost, we collect the training wall-clock time, which is 21% $\sim$ 25% less than training from scratch as the original Transformer model and more than 50$\times$ less than the previous ShiftAddNet [70] did on GPUs. For example, training the PVTv1-Tiny model from scratch takes 62 hours while our finetuning only needs 46 hours.
@@ 5.3 ShiftAddViT over SOTA Baselines on 3D Tasks
We also extend our methods to 3D NVS tasks and built ShiftAddViT on top of Transformer-based GNT models [53]. We test and compare the PSNR, SSIM, LPIPS, latency, and energy of both ShiftAddViT and baselines, i.e., vanilla NeRF [41] and GNT [53], on the LLFF dataset across eight scenes. As shown in Tab. 5, our ShiftAddViT consistently leads to better accuracy-efficiency tradeoffs, achieving 22.3%/50.4% latency reductions and 20.8%/54.3% energy savings under comparable or even better generation quality ($\uparrow$0.55$\downarrow$0.19 averaged PSNR across eight scenes), as compared to NeRF and GNT baselines, respectively. Note that GNT costs more than NeRF because of an increased number of layers. In particular, ShiftAddViT even achieves better generation quality (up to $\uparrow$0.88 PSNR) than GNT on some representation scenes, e.g., Orchid and Flower, as highlighted in Tab. 5, where the first, second, and third ranking performance are noted with corresponding colors. The full results of eight scenes are supplied in the Appendix C. Note that we only finetune ShiftAddViT for 140K steps on top of pre-trained GNT models instead of training up to 840K steps from scratch.
@@ 5.4 Ablation Studies of ShiftAddViT
     Methods Linear Attn Add Shift MoE         PVTv2-B1 [61]         PVTv2-B2 [61]
@@
PVT [61] 78.70 6.75 1512 82.00 12.20 921
Ecoformer [34] 78.38 8.02 874 81.28 15.43 483
ShiftAddViT (with KSH [34] or Quant. [27] to binarize Q/K) 78.80 6.36 1373 81.40 9.21 796
78.70 1.70 923 81.31 3.18 518
✓ (Attn) ✓ (MLPs) 78.20 2.46/1.46 N/A 81.06 4.17/3.22 N/A
✓ (Both) 78.38 2.64/1.48 N/A 81.32 4.83/3.35 N/A
@@
78.70 1.57 942 81.33 2.82 553
✓ (Both) 77.55 1.54 1021 79.97 2.93 600
✓ (Both) 78.49 2.49/1.35 N/A 81.01 4.55/3.16 N/A
Performance Breakdown Analysis. To investigate how each of the proposed techniques contributes to the final performance, we conduct ablation studies among various kinds of ShiftAddViT variants on PVTv1-T, PVTv2-B0/B1/B2 models to gradually break down and show the advantage of each component. As shown in Tab. 4 and 6, we have three general observations: (1) ShiftAddViT is robust to the binarization method of $\mathbf{Q/K}$, e.g., it is compatible with either KSH [34] or vanilla binarization [27], achieving on average 3.03$\times$ latency reductions than original PVT under comparable accuracies (±$\pm 0.5\%$), and both 3.29$\times$/1.32$\times$ latency/energy reductions and 0.04% $\sim$ 0.20% accuracy improvements over Ecoformer [34]; (2) vanilla binarization methods work better than KSH in our ShiftAddViT framework in terms of both accuracy ($\downarrow$0.05$\sim$ $\uparrow$0.21%) and efficiency (on average $\downarrow$5.9latency reductions and $\uparrow$5.8throughput boosts). Moreover, KSH requires that $\mathbf{Q}$ and $\mathbf{K}$ are identical while vanilla binarization [27] does not have such a limitation. That explains why applying vanilla quantization to the linear attention layers results in an accuracy improvement of on average 0.15% as compared to using KSH. Also note that our adopting linear attention following [73] works better ($\uparrow$0.29%) than PVT [60, 61]; (3) Replacing linear layers in attention modules or MLPs with Shift layers lead to an accuracy drop while our proposed MoE can help to compensate for it. Specifically, adopting Shift layers leads to on average 1.67% accuracy drop while MoE instead improves on average 1.37% accuracy. However, MoE hurts the efficiency of both the PVT baseline and our ShiftAddViT without customized system acceleration, e.g., PVT+MoE increases 94% latency and reduces 51% throughput than PVT on GPUs, due to limited parallelism supports, especially for TVM. We report the modularized latency by separately optimizing each expert to demonstrate the potential of MoE’s double-winning accuracy ($\uparrow$0.94$\sim$ $\uparrow$2.02%) and efficiency ($\downarrow$15.4$\sim$ $\uparrow$42.7%).
@@
Figure 4: MLP/Linear latency speedups using shifts, where inputs are of shape $B\times K\times M$, weights are of shape $K\times N$. All dimensions are set w.r.t. the real dimensions in PVTs [60].
@@
Speedups of Shifts and Adds. Apart from the overall comparison and energy reduction profiling, we also test the GPU speedups of our customized Shift and Add kernels, as shown in Fig. 5 and 5, respectively. We can see that both customized kernels achieve faster speeds than PyTorch and TVM baselines. For example, our MatAdds achieve on average 7.54$\times$/1.51$\times$ speedups than PyTorch and TVM MatMuls, respectively. Our MatShifts achieve on average 2.35$\times$/3.07$\times$/1.16$\times$ speedups than PyTorch FakeShifts [17], TVM FakeShifts, and TVM MatMuls, respectively. More comparisons and speedup analysis under larger batch sizes are supplied in the Appendix A. Note that in those comparisons, we use the default einsum operator for the PyTorch MatMuls, floating-point multiplication with power-of-twos for FakeShift, and implement the MatAdds and MatShifts using TVM [10] ourselves. We compare those customized operators with their multiplication counterparts in the same setting for fair comparisons. The speedup of MatAdds is because we replace multiplication-and-accumulation with accumulation only. The speedup of MatShifts is mainly attributed to the bit reductions (INT32 and INT8 for inputs and shift signs/weights) and thus data movement reductions instead of computations which are almost fully hidden behind data movements.